OpenStudy (anonymous):

Did I do this proof correctly? cos(-Θ)/1-sin(-Θ) = secΘ+tanΘ cosΘ/1-sinΘ cosΘ/1 - cosΘ/sinΘ = secΘ+tanΘ

OpenStudy (amistre64):

you need better use of paranthesis/brackets

OpenStudy (amistre64):

andno

OpenStudy (amistre64):

use t, or x , or a instead of the theta, makes typing easier

OpenStudy (anonymous):

oh i see okay

OpenStudy (anonymous):

(cos a)/(1-sin a) (cos a/1)-(cos a/sin a)

OpenStudy (amistre64):

fractions .... \[\frac{a}{x+y}\ne\frac ax + \frac ay \]

OpenStudy (anonymous):

ohh but if it was (x+y)/a then it would equal (a/x)+(a/y) ?

OpenStudy (amistre64):

almost ... if it was (x+y)/a then it would equal (x/a)+(y/a) ^^^^^^^^^

OpenStudy (anonymous):

Okay thank you! Just wanted to check

OpenStudy (amistre64):

yep